OpenStudy (anonymous):

Solve: 3(x + 1) + 6 = 33

hartnn (hartnn):

first subtract 6 from both sides, what u get ?

OpenStudy (anonymous):

3(x + 1) + 6 = 33 3x + 3 + 6 = 33 3x + 9 = 33 - 9 - 9 3x/3 = 24/3 x = 8 3(8 + 1) + 6= 24 + 3 + 6 = 24 + 9 = 33